OpenStudy (anonymous):

f a faucet drips at the rate of a drop every 2 seconds and it takes 575 drops to fill a 100-ml bottle, how many liters of water are wasted in a week? (Round your final answer to the nearest liter)

OpenStudy (anonymous):

No. of days in a week = 7 No. of hrs in a week = 24 * 7 = 168 hrs No. of min. in a week = 168 * 60 = 10080 mins No. of seconds in a week = 10080 * 60 = 604800 seconds 604800/2 = 302400 302400/575 = 526 526 * 100 = 52600 ml So amt. of water wasted = 52600 ml = 52.6 l(this is an approximation - the actual value of 302400/575 is 525.9)

OpenStudy (anonymous):

Are you there Monkey98?

OpenStudy (anonymous):

Yes, thank you for your help!